OpenStudy (anonymous):

x^2-4 (x-2) ------ / ------ x-8 x-3

OpenStudy (anonymous):

\[\frac{x^2-4}{x-8}\div\frac{x-2}{x-3}=\frac{x^2-4}{x-8}\times \frac{x-3}{x-2}\] \[=\frac{(x-2)(x+2)(x-3)}{(x-8)(x-2)}\] then cancel the one common factor top and bottom

OpenStudy (anonymous):

thank you ! (:

OpenStudy (anonymous):

yw